OpenStudy (anonymous):

Simplify the following expression. 2^-8*2^4 A. 16 B.1/256 C. 1/16 D. 256

OpenStudy (trojanpoem):

You sum powers: 2^-8 + 4 = 2^-4 = negative power = 1/ 1/2^4 = 1/16

pooja195 (pooja195):

\[\huge~\rm~\bf~(2^{−8})(2^4) \] Since it's -8 we need to make a fraction \[~\rm~ \frac{ 1 }{ 2*2*2*2*2*2*2*2 } \times (2*2*2*2)\]

OpenStudy (anonymous):

can u hlp with more

pooja195 (pooja195):

I didn't even help you with dis one yet >_>

pooja195 (pooja195):

Cancel out as many 2's as you can how many 2's are we left with?

OpenStudy (anonymous):

1/256x (2x 2 x2 x2)

OpenStudy (anonymous):

1/16

pooja195 (pooja195):

No we need to cancel out all the 2's we can \[\huge~\rm~\bf~(2^{−8})(2^4)\] \[~\rm~ \frac{ 1 }{\color{red}{ 2*2*2*2*2*2*2*2} } \times (\color{green}{2*2*2*2})\]

pooja195 (pooja195):

we have 8 2's that are red and 4 2's that are green 8-4= ?

OpenStudy (anonymous):

ohh

pooja195 (pooja195):

8-4=?

OpenStudy (anonymous):

4

pooja195 (pooja195):

Good 1 remains the same \[\huge~\rm~\bf~\frac{ 1 }{ 2*2*2*2 }\] so we have \[\huge~\rm~\bf~\color{red}{\frac{ 1 }{ 2^4 }}\] \[\huge~\rm~\bf~\color{purple}{\frac{ 1 }{ 2^4=? }}\]

OpenStudy (anonymous):

1/16

pooja195 (pooja195):

Well done :-)

OpenStudy (anonymous):

Simplify the following expression. 5^-2/5^-4 A. 15,625 B. 1/15,625 C. 25 D. 1/25

OpenStudy (anonymous):

C ?

pooja195 (pooja195):

\[\huge~\rm~ \frac{ a^b }{ a^c } \rightarrow a^{b-c }\]

pooja195 (pooja195):

So, \[\huge~\rm~\bf~ \frac{ 5^{-2} }{ 5^{-4} }\] You would subtract -2-4 \[\huge~\rm~\bf~\color{purple}{5^{−2−(−4)}}\]

OpenStudy (anonymous):

-(- doesnt that mean add

pooja195 (pooja195):

Yes

OpenStudy (anonymous):

5^-2+4

pooja195 (pooja195):

Yes so add 4

OpenStudy (anonymous):

5^2=25

OpenStudy (anonymous):

thx

pooja195 (pooja195):

Well done yw ^_^

OpenStudy (anonymous):

Simplify the following expression. 2^5x 2^-7

OpenStudy (anonymous):

(25)(2−7) =25*2−7 =(2*2*2*2*2)*( 1 2*2*2*2*2*2*2 ) = 1 2*2 = 1 22 = 1 4 (Decimal: 0.25)

pooja195 (pooja195):

Nice calculator answer.

OpenStudy (anonymous):

thx for your hlp

pooja195 (pooja195):

yw :P